OpenStudy (anonymous):

how do i find out if these two lines are perpendicular? please show me how to work the problem: -4x+2y=12, x+2y=6

OpenStudy (aroub):

Do you know how to find the slope ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well maybe not

OpenStudy (aroub):

Okay, great! What are the slopes?

OpenStudy (aroub):

lol

OpenStudy (aroub):

Fine I'll tell you how :)

OpenStudy (anonymous):

i get the y =mx+b? or is there another method that would make this easier?

OpenStudy (anonymous):

thanks

OpenStudy (aroub):

No you got it! y=mx+b where 'm' is the slope and 'b' is the y-intercept.

OpenStudy (anonymous):

ok, then what do i do to the problem? first i get them in the correct format? and then?

OpenStudy (aroub):

I'll do the first one and you do the second one. -4x+2y=12 As you said you return it to the form y=mx+b 2y=4x+12 y=2x+6 And right we said that y=mx+b where 'm' is the slope ? What do we have instead of m here ? 2 right ? For the first one the slope is 2. Get it ?

OpenStudy (anonymous):

yes, the second is y=-1/2x+3, so then what

OpenStudy (aroub):

Yep! What is the slope ?

OpenStudy (anonymous):

-1/2, now how do i know if they are perpendicular?

OpenStudy (aroub):

Correct, you know they are perpendicular when you multiply the slope together you get -1 -1/2*2=-1 So yes they are perpendicular.

OpenStudy (anonymous):

so if they equal -1 when multiplied they are perp? then are they parallel if they slop is similar and the y intercept is different?

OpenStudy (aroub):

Yes! Don't worry about the y-intercept.

OpenStudy (anonymous):

why not worry about the y intercept?

OpenStudy (aroub):

Because you won't use it to know if they are perpendicular or parallel.

OpenStudy (anonymous):

so is a line parallel if it has the same slope always?

OpenStudy (aroub):

Yes!

OpenStudy (anonymous):

oh, thanks

OpenStudy (aroub):

Anytime! =D